# Rng: Relightable Neural Gaussians

### v4

@@ Abstract
3D Gaussian Splatting (3DGS) has shown impressive results for the novel view synthesis task, where lighting is assumed to be fixed. However, creating relightable 3D assets, especially for objects with ill-defined shapes (fur, fabric, etc.), remains a challenging task. The decomposition between light, geometry, and material is ambiguous, especially if either smooth surface assumptions or surface-based analytical shading models do not apply. We propose Relightable Neural Gaussians (Rng), a novel 3DGS-based framework that enables the relighting of objects with both hard surfaces or soft boundaries, while avoiding assumptions on the shading model. We condition the radiance at each point on both view and light directions. We also introduce a shadow cue, as well as a depth refinement network to improve shadow accuracy. Finally, we propose a hybrid forward-deferred fitting strategy to balance geometry and appearance quality. Our method achieves significantly faster training (1.3 hours) and rendering (60 frames per second) compared to a prior method based on neural radiance fields and produces higher-quality shadows than a concurrent 3DGS-based method.
@@ 3.3 Shadow cue
An intuitive and naive proxy for the depth is the weighted sum of the depth and opacity of each Gaussian,
@@
where $z_{i}$ is the depth value of $i^{\textrm{th}}$ Gaussian on the camera ray and $\alpha_{i}$ is its opacity. However, sometimes the weighted sum is incorrect, leading to wrongly located shading points and consequently mismatching shadow cues. We discuss and showcase this situation in the supplementary. To address this issue, we propose a depth refinement network to correct the shading point locations by learning a scaling factor, as shown in Fig. 2. We assume the depth correction is linear for each pixel and dependent on view directions $\omega_{o}$. Therefore, the refined depth value is obtained by
@@
where $V$ is the shadow cue. In practice, we use the same resolution as the camera for the shadow camera, and we apply a clamping between zero and the scene units to the shadow cue map for the stability of training.
@@ 4 Results
Scene GS-IR[18] Relightable 3D Gaussian[9] Ours
@@ 4.2 Experiment setups
We select four representative NeRF-/GS-based relighting methods for comparison. For validating the relighting quality under point lights, we compare with NRHints [27] and GS3 [2]. Furthermore, we compare with GS-IR [18] and Relightable 3D Gaussian [9] to validate the relighting under environment lights. Note that NRHints and GS3 also describes an additional optimization on camera poses to compensate for the calibration precision of NRHints datasets, but we disable it for a fair comparison.
@@ 4.3 Quality validation
In our supplementary materials, we provide additional validation and visualization of our learned geometry and shadows and the relighting quality with moving lights/views. We also validate the power of neural appearance by comparing it to vanilla 3DGS. Please refer to them for more details.
@@ 4.5 Discussion and limitations
Although we deploy the shadow cue to help the network predict better shadow appearances, the shadow quality is still limited by the geometry reconstruction precision. A perfect geometry reconstruction for soft-boundary objects is not trivial. Therefore, we suffer from this disadvantage like most GS-based approaches.

### v5

@@ Abstract
3D Gaussian Splatting (3DGS) has shown impressive results for the novel view synthesis task, where lighting is assumed to be fixed. However, creating relightable 3D assets, especially for objects with ill-defined shapes (fur, fabric, etc.), remains a challenging task. The decomposition between light, geometry, and material is ambiguous, especially if either smooth surface assumptions or surface-based analytical shading models do not apply. We propose Relightable Neural Gaussians (RNG), a novel 3DGS-based framework that enables the relighting of objects with both hard surfaces or soft boundaries, while avoiding assumptions on the shading model. We condition the radiance at each point on both view and light directions. We also introduce a shadow cue, as well as a depth refinement network to improve shadow accuracy. Finally, we propose a hybrid forward-deferred fitting strategy to balance geometry and appearance quality. Our method achieves significantly faster training (1.3 hours) and rendering (60 frames per second) compared to a prior method based on neural radiance fields and produces higher-quality shadows than a concurrent 3DGS-based method. Project page: whois-jiahui.fun/project_pages/RNG.
@@ 3.3 Shadow cue
An intuitive and naive proxy for the depth is the weighted sum of the depth and Gaussian responses along the ray,
@@
where $z_{i}$ is the depth value of $i^{\textrm{th}}$ Gaussian on the camera ray and $\alpha_{i}$ is the ray response at each intersections of 3D Gaussians. Usually, this proxy is normalized, thus the background leakage at semi-transparent areas can be reduced. However, sometimes the weighted sum is incorrect, leading to wrongly located shading points and consequently mismatching shadow cues. We discuss and showcase this situation in the supplementary. To address this issue, we propose a depth refinement network to correct the shading point locations by learning a scaling factor, as shown in Fig. 2. We assume the depth correction is linear for each pixel and dependent on view directions $\omega_{o}$. Therefore, the refined depth value is obtained by
@@
where $V=|PQ|$ is the shadow cue. In practice, we use the same resolution as the camera for the shadow camera, and we apply a clamping between zero and the scene units to the shadow cue map for the stability of training.
@@ 3.4 Hybrid optimization
Scene GS-IR[18] RelightableGS[9] Ours
@@ 4.2 Experiment setups
We select four representative NeRF-/GS-based relighting methods for comparison. For validating the relighting quality under point lights, we compare with NRHints [27] and GS3 [2]. Furthermore, we compare with GS-IR [18] and Relightable 3D Gaussian [9] to validate the relighting under environment lights.
@@ 4.3 Quality validation
We also provide additional validation and visualization of our learned geometry and shadows, the relighting quality, and the power of neural appearance in our supplementary. Please refer to them for more details.
@@ 4.5 Discussion and limitations
Although we deploy the shadow cue to help the network predict better shadow appearances and present the depth refinement network to compensate for this inaccuracy, the shadow quality is still limited by the geometry reconstruction precision, since a perfect geometry reconstruction for soft-boundary objects is not trivial. Therefore, we suffer from this disadvantage like most GS-based approaches.
@@ 5 Conclusion
Acknowledgements
We thank the reviewers for their valuable comments. This work is supported by the National Science Fund of China under Grant Nos. U24A20330, 62361166670, and 62172220.
@@ 6 Additional validation
In Fig. 13, we demonstrate the effectiveness of our shadow cue by showcasing an example where the shadow quality dominates. We observe a significant quality improvement in the shadow by adding shadow cues into our model.
@@
Figure 13: The quality difference at shadow regions with/without the shadow cues. With the shadow cues applied, the network can further improve the quality of dark pixels, as well as provide a clearer boundary for the shadow region.